Luke has $5 more than Sam. Together they have $73.00. How much does each have?
Answer by emargo19(101)

Let Sam have x amount
Therefore Luke has x+5
x+x+5=73
2x=73-5
2x=68
x=34
Sam has $34.00
Luke has x+5
=34+5
=39
So Luke has $39.00